OpenStudy (anonymous):

can someone help me find the derivative of this problem? f(x)= ln (sqrt of 1+sin^2x)

OpenStudy (amistre64):

chain rule practice !!!

OpenStudy (loser66):

Let me help you write it in neat: \(ln\sqrt{1+sin^2x}\)

OpenStudy (amistre64):

what is the derivative of ln(u) ?

OpenStudy (anonymous):

1/u

OpenStudy (amistre64):

almost: u is a function, so the chain rule applies: implicitly we have: ln(u) derives to u'/u but u = sqrt(v), therefore what is u'?

OpenStudy (amistre64):

edit: u' = v' / 2sqrt(v) but v = 1 + sin^2(x), so what is v'? ln(u(v(x))) = 1/u * 1/2sqrt(v) * v' or: ln(u(v(x))) = v'/ 2usqrt(v)

OpenStudy (anonymous):

uhm well v'= 1/2(1+sin^2x)^-1/2 (2)(sinx)(cosx)

OpenStudy (amistre64):

its a chain rule practice problem ... just gonna be messy messy messy

OpenStudy (amistre64):

\[\frac{d}{dx}ln(\sqrt{1+sin^2x})\] \[\frac{\frac d{dx}(\sqrt{1+sin^2x})}{\sqrt{1+sin^2x}}\] \[\frac{\frac d{dx}(1+sin^2x)}{2\sqrt{1+sin^2x}\sqrt{1+sin^2x}}\] \[\frac{\frac d{dx}(1)+\frac d{dx}(sin^2x)}{2(1+sin^2x)}\] \[\frac{2sin(x)\frac d{dx}(sinx)}{2(1+sin^2x)}\] \[\frac{2sin(x)cos(x)}{2(1+sin^2x)}\]

OpenStudy (amistre64):

spose we can cancel the 2s and such but i think thats about it

OpenStudy (anonymous):

okay wait but why divide by sqrt of 1+ sin^2x

OpenStudy (anonymous):

no actually how did you get sqrt of 1+sin^2x in the numerator?